OpenStudy (anonymous):

25xy^4+60xy^3divided by 5xy

OpenStudy (mertsj):

\[\frac{25xy^4+60xy^3}{5xy}=\frac{25xy^4}{5xy}+\frac{60xy^3}{5xy}\]

OpenStudy (mertsj):

Does that help?

OpenStudy (phi):

Let's concentrate on the first part \[ \frac{25xy^4}{5xy} \] do you know what 25/5 is ?